OpenStudy (anonymous):

Which expression is equivalent to 8^12/8^3 ? A.)1^9 B.) 8^4 C.) 8^9 D.) 5^12 Show your work

OpenStudy (bibby):

When dividing exponents with a common base you can subtract the powers like so":\[\frac{ a^m }{ a^n } = a^{m-n}\]

OpenStudy (mathstudent55):

\(\dfrac{a^m}{a^n} = a^{m - n} \)

OpenStudy (the_fizicx99):

Go bib!!